How do you answer questions like this one : 2x+20=100

Answer by kaykay714(2)

Okay, you want to get x alone. First, you subtract 20 from the 20 and the 100. But you do not do anything to the 2x. You will deal with the 2x soon. So now you have
2x+20=100
.......-20=-20
________
2x + 0 = 80
So now you don't need the + 0 in there any more. So you change it to; 2x = 80. You still want x alone. So you divide 2x and 80 by 2x. So it looks like this;
2x = 80
______
2x......2x
x = 40
If it was 2x-20=100, then you do the same exact thing, accept you would add the 20 like you did when you subtracted it.
(dont mind the .'s, i just put that in there so the problem would look right as if they weren't there).
~kelly~